# Scalable Multi-view Clustering via Explicit Kernel Features Maps

### v1

@@ 4 Proposed Methodology
The problem of applying spectral clustering via the proper decomposition of this matrix would result in a computational complexity of $\mathcal{O}(n^{2}k)$ and a spatial complexity of $\mathcal{O}(n^{2})$. This means that we would lose the efficiency argued by [5] and thus lose the advantage of using such an approach in the first place. Therefore, we propose solving this problem via the summation property of kernels.
@@ 4.1 Weighing Views with respect to their Clustering Structure
We additionally introduce a weighting scheme on the view-specific terms in problem 6, since the different views should not always have the same contribution level towards the consensus affinity matrix. Here, we evaluate the importance of each view using the affinity matrix corresponding to that view. Specifically, we consider the following problem:
@@
such that $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$
$\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
@@
$\text{such that}\quad\forall w\in[V]\qquad\ \ \mathbf{W}_{v}=\Phi(\mathbf{U}_{w})\Phi(\mathbf{U}_{w})^{\top}$
$\qquad\qquad\qquad\qquad\qquad\quad\mathbf{G}_{w}\in\{0,1\}^{n\times k},\quad\mathbf{G}_{w}\mathbf{1}=\mathbf{1}$
where $T$ is a temperature parameter and $\mathbf{G}_{1},\ldots,\mathbf{G}_{V}$ are different view-specific partition matrices that for spectral clustering we solve by applying singular value decomposition on $\mathbf{D}_{w}^{-\frac{1}{2}}\Phi(\mathbf{U}_{w})$ and then applying a clustering algorithm on the resulting left singular vectors and repeating the process for each $v\in[V]$.
4.2 Multi-view Scalability via Kernel Summation
Property 1.
This means that matrix 7 is also a kernel and, consequently, has a decomposition of the form $\mathbf{B}\mathbf{B}^{\top}$ for some matrix $\mathbf{B}$ (this is property of kernel matrices). Additionally, we have the obvious fact:
Property 2.
When $k_{a}$ and $k_{b}$ are nonnegative kernels, i.e. $\forall x,y$ $k_{a}(x,y)\geq 0$ and $k_{b}(x,y)\geq 0$, then $k(x,y)=k_{a}(x,y)+k_{b}(x,y)$ is also a nonnegative kernel.
This means that every entry of 7 is nonnegative, making it a valid affinity matrix. To solve the efficiency issue of combining the different view-specific affinity matrices we therefore need a way of obtaining a vector realization $\mathbf{B}$ for the decomposition of 7. Once we have $\mathbf{B}$, we can obtain a spectral clustering via singular value decomposition of $\mathbf{B}$ instead of eigendecomposition of $\mathbf{B}\mathbf{B}^{\top}$. To this end we make use of the following property:
Property 3.
Using this property, we can infer that a vector realization is $\mathbf{B}=\left[\mathbf{D}_{1}^{-\frac{1}{2}}\Phi(\mathbf{U}_{1}),\ldots,\mathbf{D}_{V}^{-\frac{1}{2}}\Phi(\mathbf{U}_{V})\right]$. This decomposition means that the complexity of the approach is proportional to dimension $m$ of the feature map $\Phi$. This leads to large speedup and memory saving when $m<<n$.
@@ 4.3 Optimization
In addition to the use of properties associated with kernel feature maps, the efficiency of our approach also stems from the fact that we do not use an iterative algorithm, in contrast to many other state-of-the-art models, e.g, [17]. The algorithm consists of the following steps:
@@ 4.3.1 Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$
$\mathbf{C}_{1},\ldots,\mathbf{C}_{v-1},\mathbf{C}_{v+1},\ldots,\mathbf{C}_{V}$ and $F$ in problem 6 and reiterating for each $v\in[V]$. The resulting problem is 3, and the solution is to set $\mathbf{U}$ as the $f$ left singular vectors of $\mathbf{X}$ associated with the largest singular values. Note that to solve our approach in one step and for scalability, as opposed to [7], we consider $\mathbf{W}_{v}$ not to be a function of $\mathbf{C}_{v}$.
@@ 4.3.2 Computing $\lambda_{1},\ldots,\lambda_{V}$
Once we have obtained the singular vectors associated with each coefficient matrix $\mathbf{C}_{v}$, we compute the corresponding $\lambda_{v}$ using Formula 9.
@@ 4.3.3 Solving for $\mathbf{F}$
$\min_{\mathbf{F}}\quad$ $\sum_{v}\texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})$ (10)
such that $\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}.$
@@
The solution $\mathbf{F}$ of this problem is obtained by applying $k$-means clustering on the left singular values corresponding to the largest $k$ singular values of matrix $\oplus_{v}\ \sqrt{\lambda_{v}}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})$.
A pseudo-code for our approach is given in algorithm 1.
@@ 4.4 Complexity Analysis
    Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$ This process takes $O(vnd\log(f))$ where $f$ is the chosen rank due to the randomized singular values decomposition.
@@
    Computing $\lambda_{1},\ldots,\lambda_{V}$ The complexity corresponds to computing the projections using feature map $\Phi$. We consider the case of using a quadratic polynomial feature map whose computational complexity is $O(nf^{2})$. The computational complexity of this step is then $O(Vnf^{2}log(k))$.
@@
    Solving for $\mathbf{F}$ Solving for $\mathbf{F}$ Finally, the process of computing the concatenated matrix $\mathbf{B}$ and then applying singular value decomposition is $O(Vnf^{2}\log(k))$. Applying k-means afterwards results in $O(tVnf^{2}k)$ operations.
@@
    The effect of the choice of the non-negative Kernel It is possible to use nonnegative kernels other than the quadratic polynomial. However, in the case of infinite-dimensional kernels such as RBF and sigmoid kernels or, in the case of polynomials, kernels with large even degrees, it is necessary to use kernel feature map approximation techniques such as Nystroem [26] or polynomial sketching [18]. In these cases, the factor $f^{2}$ is replaced by the dimension of the feature map approximation, and the complexity of computing the approximation of the kernel has to be added to the total complexity of our approach.
In table 1 we report the complexity of several state-of-the-art algorithms. We see that considerably more factors must be taken into consideration than are needed when using our approach. Note that we have added the complexity of learning graph representations (using 12) to non-graph methods.
Data: Data matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, number of clusters $k$, number of components $f$, kernel feature map $\Phi$, temperature $T$
Result: Consensus partition matrix $\mathbf{F}$
for  $v\in[V]$ do
       Subtract column means from $\mathbf{X}_{v}$;
       Construct matrix $\mathbf{U}_{v}$ by stacking the $f$ first left singular vectors of $\mathbf{X}_{v}$ as its columns;
       Compute $d$ the row sums of $\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$ efficiently;
       $\mathbf{B}_{v}=\mathbf{B}_{v}*\textbf{d}^{-0.5}$;
       Construct matrix $\mathbf{Q}_{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}_{v}$ as its columns and then drop the first column;
       Compute partition matrix $\mathbf{G}_{v}$ via k-means on $\mathbf{Q}_{v}$ with $k$ clusters
Compute $\lambda_{1},...,\lambda_{V}$ using Formula 9 with temperature $T$, $\mathbf{G}_{v}$ and $\mathbf{W}_{v}=\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$ ;
$\mathbf{B}=\oplus_{v}\lambda_{v}\mathbf{B}_{v}$ ;
Compute $d$ the row sums of $\mathbf{B}\mathbf{B}^{\top}$ efficiently ;
$\mathbf{B}=\mathbf{B}*\textbf{d}^{-0.5}$ ;
Construct matrix $\mathbf{Q}\phantom{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}\phantom{v}$ as its columns and then drop the first column ;
Compute partition matrix $\mathbf{F}$ via k-means on $\mathbf{Q}$ with $k$ clusters
Algorithm 1 MvSCK
Table 1: Complexity of some of the considered approaches on network data. $m$ represents the number of anchors. $|E|$ is the number of edges in the graphs used in Formula 12. For the sake of simplicity, we consider that the dimension of the features and the number of graph edges in the different views are the same.
Method Time Complexity
@@
MvSCK $O(Vp|E|d+Vnf^{2}\log(k))$
Table 2: Characteristics of the Datasets. The imbalance is given as the ratio between the cardinalities of the most frequent and least frequent classes.
@@
Amazon Photos Amazon Computers
@@
MvSCK 75.66 ±2.2 66.56 ±5.0 70.66 ±1.3 58.65 ±2.0 66.34 ±1.5 56.23 ±0.8 57.15 ±0.8 46.27 ±2.0
Table 5: Clustering results on the large scale datasets. We allow for a maximum runtime of two hours per
run otherwise we report a timeout.
OGBN-ArXiv OGBN-Products
@@
MvAGC Out of Memory Out of Memory
MAGC Out of Memory Out of Memory
@@
MvSCK 48.29 ±1.8 28.46 ±0.9 46.99 ±0.5 39.15 ±1.2 47.14 ±1.5 16.89 ±1.3 50.80 ±0.8 32.24 ±1.3

### v2

@@ 5 Proposed approach
However, applying spectral clustering through the proper decomposition of this matrix would lead to a computational complexity of $\mathcal{O}(n^{2}k)$ and a spatial complexity of $\mathcal{O}(n^{2})$. This implies that we would lose the efficiency argued by [fettal2023scalable] and thus lose the advantage of using such an approach in the first place. Therefore, we propose addressing this problem via the summation property of kernels.
5.1 Multi-view Scalability via Kernel Summation
proper 1.
This means that matrix (8) is also a kernel and, consequently, has a decomposition of the form $\mathbf{B}\mathbf{B}^{\top}$ for some matrix $\mathbf{B}$ (this is the property of kernel matrices). Additionally, we have the obvious fact:
proper 2.
When $k_{a}$ and $k_{b}$ are non-negative kernels, i.e. $\forall x,y$ $k_{a}(x,y)\geq 0$ and $k_{b}(x,y)\geq 0$, then $k(x,y)=k_{a}(x,y)+k_{b}(x,y)$ is also a nonnegative kernel.
This means that every entry of (8) is non-negative, making it a valid affinity matrix. To solve the efficiency issue of combining the different view-specific affinity matrices, we therefore need a way of obtaining a vector realization $\mathbf{B}$ for the decomposition of (8). Once we have $\mathbf{B}$, we can obtain a spectral clustering by decomposition of singular values of $\mathbf{B}$ instead of the eigendecomposition of $\mathbf{B}\mathbf{B}^{\top}$. To this end, we make use of the following property:
proper 3.
Using this property, we find that a realization is $\mathbf{B}=\left[\mathbf{D}_{1}^{-\frac{1}{2}}\Phi(\mathbf{U}_{1}),\ldots,\mathbf{D}_{V}^{-\frac{1}{2}}\Phi(\mathbf{U}_{V})\right]$. This decomposition means that the complexity of the approach is proportional to dimension $m$ of the feature map $\Phi$. This leads to large speedup and memory saving when $m<<n$.
@@ 5.2 Weighing Views with respect to their Clustering Structure
We additionally introduce a weighting scheme on the view-specific terms in problem (7), since different views should not always have the same contribution level towards the consensus affinity matrix. Here, we evaluate the importance of each view using the affinity matrix corresponding to that view. To be more precise, the proposed objective function changes into, with the same constraints:
@@
s.t. $\mathbf{W}_{v}=\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top}\mathbf{D}_{v}^{-\frac{1}{2}},\quad\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}$
@@
s.t. $\forall v\in[V]\ \mathbf{W}_{v}=\Phi(\mathbf{U}_{v})\Phi(\mathbf{U}_{v})^{\top},\quad\mathbf{G}_{v}\in\{0,1\}^{n\times k},\mathbf{G}_{v}\mathbf{1}=\mathbf{1}$
where $T$ is a temperature parameter and $\mathbf{G}_{1},\ldots,\mathbf{G}_{V}$ are different view-specific partition matrices.
We solve this problem via spectral clustering: we apply singular value decomposition on $\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})$ and then apply a clustering algorithm on the resulting left singular vectors, repeating the process for each $v\in[V]$.
@@ 5.3 Optimization
In addition to the use of properties associated with kernel feature maps, the efficiency of our approach also stems from the fact that we do not use an iterative algorithm, in contrast to many other state-of-the-art models, e.g, [pan2021multi]. The algorithm consists of the following steps:
@@ 5.3.1 Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$
$\mathbf{C}_{1},\ldots,\mathbf{C}_{v-1},\mathbf{C}_{v+1},\ldots,\mathbf{C}_{V}$ and $\mathbf{F}$ in problem (7) and reiterating for each $v\in[V]$. The resulting problem is (4), and the solution is to set $\mathbf{U}$ as the $f$ left singular vectors of $\mathbf{X}$ associated with the largest singular values. Note that to solve our approach in one step and for scalability, we consider $\mathbf{W}_{v}$ to be not a function of $\mathbf{C}_{v}$.
@@ 5.3.2 Computing $\lambda_{1},\ldots,\lambda_{V}$
Once we have obtained the singular vectors associated with each coefficient matrix $\mathbf{C}_{v}$, we compute the corresponding $\lambda_{v}$ using Formula (LABEL:eq:lambda).
@@ 5.3.3 Solving for $\mathbf{F}$
$\min_{\mathbf{F}}$ $\sum_{v}\texttt{Tr}(\mathbf{F}^{\top}(\mathbf{I}-\mathbf{W}_{v})\mathbf{F})\text{such that}\quad\mathbf{F}\in\{0,1\}^{n\times k},\quad\mathbf{F}\mathbf{1}=\mathbf{1}.$ (11)
@@
The solution $\mathbf{F}$ of this problem is obtained by applying $k$-means clustering on the left singular values corresponding to the largest $k$ singular values of matrix $\oplus_{v}\ \sqrt{\lambda_{v}}\mathbf{D}_{v}^{-\frac{1}{2}}\Phi(\mathbf{U}_{v})$. A pseudo-code for our approach is given in Algorithm 1.
0:  Data matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, number of clusters $k$, number of components $f$, kernel feature map $\Phi$, temperature $T$
0:  Consensus partition matrix $\mathbf{F}$
1:for $v\in[V]$ do
2:  Subtract column means from $\mathbf{X}_{v}$
3:  Construct matrix $\mathbf{U}_{v}$ by stacking the $f$ first left singular vectors of $\mathbf{X}_{v}$ as its columns
4:  $\mathbf{B}_{v}=\Phi(\mathbf{U}_{v})$
5:  Compute $d$ the row sums of $\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$ efficiently
[…]
7:  Construct matrix $\mathbf{Q}_{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}_{v}$ as its columns and then drop the first column
8:  Compute partition matrix $\mathbf{G}_{v}$ via k-means on $\mathbf{Q}_{v}$ with $k$ clusters
9:end for
10: Compute $\lambda_{1},...,\lambda_{V}$ using Formula (LABEL:eq:lambda) with temperature $T$, $\mathbf{G}_{v}$ and $\mathbf{W}_{v}=\mathbf{B}_{v}\mathbf{B}_{v}^{\top}$
[…]
12: Compute $d$ the row sums of $\mathbf{B}\mathbf{B}^{\top}$ efficiently
[…]
14: Construct matrix $\mathbf{Q}\phantom{v}$ by stacking the $f+1$ first left singular vectors of $\mathbf{B}\phantom{v}$ as its columns and then drop the first column
15: Compute partition matrix $\mathbf{F}$ via k-means on $\mathbf{Q}$ with $k$ clusters
Algorithm 1 Pseudocode for MvSCK
@@ 5.4 Complexity Analysis
    Solving for $\mathbf{C}_{1},\ldots,\mathbf{C}_{V}$
    This process takes $O(vnd\log(f))$ where $f$ is the chosen rank due to the randomized singular values decomposition.
@@
    Computing $\lambda_{1},\ldots,\lambda_{V}$
    The complexity corresponds to computing the projections using feature map $\Phi$. We consider the case of using a quadratic polynomial feature map whose computational complexity is $O(nf^{2})$. The computational complexity of this step is then $O(Vnf^{2}log(k))$.
@@
    Solving for $\mathbf{F}$
    Finally, the process of computing the concatenated matrix $\mathbf{B}$ and then applying singular value decomposition is $O(Vnf^{2}\log(k))$. Applying k-means afterwards results in $O(tVnf^{2}k)$ operations where $t$ denotes the number of iterations.
@@
    The effect of the choice of the non-negative Kernel
    It is possible to use nonnegative kernels other than the quadratic polynomial. However, in the case of infinite-dimensional kernels such as RBF and sigmoid kernels or, in the case of polynomials, kernels with large even degrees, it is necessary to use kernel feature map approximation techniques such as Nystroem [williams2000using] or polynomial sketching [pham2013fast]. In these cases, the factor $f^{2}$ is replaced by the dimension of the feature map approximation, and the complexity of computing the approximation of the kernel has to be added to the total complexity of our approach.
In table 1 we report the complexity of several state-of-the-art algorithms. We see that considerably more factors must be taken into account than are needed when using our approach. Note that we have added the complexity of learning graph representations (using (13)) to non-graph methods.
Table 1: Complexity of some of the considered approaches on network data. $m$ represents the number of anchors. $|E|$ is the number of edges in the graphs used in Formula (13). For the sake of simplicity, we consider that the dimension of the features and the number of graph edges in the different views are the same.
Method
Time Complexity
@@
MvSCK
$O(Vp|E|d+Vnf^{2}\log(k))$
6 Experiments
  • How does MvSCK compare with state-of-the-art in terms of:
    • RQ1
      : Clustering performance?
    • RQ2
      : Computational efficiency?
  • RQ3:
    Does our approach consistently outperform the state-of-the-art ?
  • RQ4:
    How do the hyper-paramters affect MvSCK?
  • RQ5:
    Is the view-importance parameter necessary?
Possible future research questions which could be explored in the future include extension to co-clustering [fettal2022subspace, fettal2023boosting], incomplete-views [xu2015multi, jiang2024deep] and streaming data [shamsinejad2025providing, shamsinejad2024anonymizing], etc.
6.1 Datasets
We evaluate our method on six benchmark datasets with different degrees of overlap: two small scale datasets, namely ACM and DBLP [fan2020one2multi]; two medium datasets, namely Photos and Computers [shchur2018pitfalls]; and two large datasets ArXivand Products [hu2020open]. The characteristics of the datasets are reported in Table 2. The six datasets are described briefly below:
  • ACM
    is a network dataset of research papers, where the associated node features are the papers’ keywords in bag-of-words format. The “Co-Subject” and “Co-Author” graphs respectively connect papers with the same theme or the same author.
  • DBLP
    is a network dataset of authors with bag-of-words node features representing keywords. “Co-Author”, “Co-Conference”, and “Co-Term” links indicate that the authors concerned have collaborated on papers, published papers at the same conference, or published papers with the same terms.
  • Photos & Computers
    are segments of the Amazon co-purchase network dataset. The nodes represent goods, and their attribute features are represented by bag-of-words of product reviews. The edges indicate that two goods are purchased together. The second set of attributes represents the word co-occurrence matrix.
  • OGBN-ArXiv
    represents a citation network encompassing a set of computer science arXiv papers. Each paper comes with a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract. We add a second $k$-nearest neighbors (k-NN) graph with 10 neighbors and added self-connections.
  • Products
    is a graph representing an Amazon product co-purchase network. An edge between two products signifies that those products are purchased together. The node features are created by extracting bag-of-words features from the product descriptions and reducing their dimensionality to 100 via PCA. Similarly, to OGBN-ArXiv, we add a second k-NN graph.
Table 2: Characteristics of the Datasets. For example, ArXivrepresents a citation network consisting of a set of computer science arXiv papers. Each paper is associated with a 128-dimensional feature vector obtained by averaging the embeddings of words in its title and abstract. We also add a second $k$-nearest neighbors (k-NN) graph with 10 neighbors and include self-connections. The imbalance is the ratio between the most frequent and least frequent classes.
@@ 6.1 Datasets
6.2 Experimental Setup
6.2.1 Experimental setting
For the purposes of comparison, we validate our approach against the state-of-the-art approaches discussed in the Related Work section above, with reference to clustering accuracy (CA), clustering F1-score (CF1) [powers2020evaluation], adjusted rand index (ARI) [steinley2004properties], and normalized mutual information (NMI) [strehl2002cluster]. We use the official implementations of the compared approaches. We choose parameters according to the authors’ recommendations. When these recommendations are not available, we set parameters similarly to how they were set on datasets of similar dimensions. To make comparisons fair, we remove any supervised signals when running experiments in the original code; these would include, for example, the validation of runs based on supervised metrics such as accuracy or setting random seeds. In our method, we employ the quadratic kernel, which was also utilized in [fettal2023scalable] and demonstrated to be effective by the authors. The number of components parameter $f$ was set equal to the number of clusters in all datasets, with the exception of Products, where it was set to 10 due to its large size. Note that all metrics reported are averaged over five different runs. All experiments were performed on a laptop equipped with a 12th Gen Intel(R) Core(TM) i9-12950HX (24 CPU threads) 2.3 GHz processor, a 12GB GeForce RTX 3080 Ti GPU, and 64GB of RAM.
6.2.2 Learning Node Representations
To perform clustering on network datasets using generic approaches such as ours in addition to LMVSC, SMVSC, FDAGF and OMSC, we first need to learn node embeddings. Given a multi-view attributed network consisting of adjacency matrices $\mathbf{A}_{1},\ldots,\mathbf{A}_{V}$ and feature matrices $\mathbf{X}_{1},\ldots,\mathbf{X}_{V}$, we learn embeddings via Laplacian smoothing. This is a well-known technique that has been used in works such as [pan2021multi, fettal2023simultaneous] to handle graph data. It involves consecutive neighborhood averaging steps:
where $p$ is the propagation order. Specifically, in our experiments, we set the propagation order as follows: 2 for ACM, 2 for DBLP, 20 for Photos, 40 for Computers, 60 for OGBN-ArXiv, and 100 for Products. This step can be seen as a preprocessing step that is independent of the approaches themselves. It is a methodology that has emerged from work done on simplifying graph convolutional networks [wu2019simplifying].
6.3 Clustering Results (RQ1)
The clustering outcomes for small, medium, and large graph datasets are detailed in Tables 3, 4, and 5 respectively. A comprehensive analysis of these tables reveals the superior performance of our approach relative to other methodologies across a majority of datasets and evaluation metrics. Remarkably, our method exhibits superiority even when compared to specialized techniques designed explicitly for graph datasets. Noteworthy observations include LMVSC’s competitiveness with our approach on the Computers dataset, where both methods achieve comparable results. Basically, while LMVSC outperforms our approach in terms of F1 and ARI metrics, our method outperforms it in accuracy and NMI metrics. This comparison underscores the strengths of our approach, showcasing its robust performance across diverse datasets and metrics. We observed some disparities between the results reported in the original papers and those obtained by running the code on the same datasets. Additionally, many approaches struggle to scale to challenging datasets like ArXivand Products. Specifically, OMSC fails to retrieve the correct number of clusters in Photos and Computers, impacting its ability to compute CF1 scores.
@@ 6.3 Clustering Results (RQ1)
Photos
Computers
@@
MvSCK
75.66 ±2.2
66.56 ±5.0
70.66 ±1.3
58.65 ±2.0
66.34 ±1.5
56.23 ±0.8
57.15 ±0.8
46.27 ±2.0
Table 5: Clustering results on the large scale datasets. We allow for a maximum runtime of two hours per run.
OGBN-ArXiv
Products
@@
MvAGC OOM OOM
MAGC OOM OOM
@@
MvSCK
48.29 ±1.8
28.46 ±0.9
46.99 ±0.5
39.15 ±1.2
47.14 ±1.5
16.89 ±1.3
50.80 ±0.8
32.24 ±1.3
6.4 Running Times (RQ2)
The detailed breakdown of running times for different models across all datasets is presented in Table 6, where our approach emerges as the leader with markedly shorter execution times over small, medium, and large graph datasets. The notable efficiency of our method becomes especially evident when compared to alternative approaches. Notably, for the challenging ArXivand Products datasets, our approach not only maintains its speed advantage but also adheres to the strict three-hour time limit we set. In contrast, several competing models exceed this limit, highlighting the computational efficiency of our proposed method in the face of diverse datasets.